Question 18 Not yet answered demand function is P 70 - 2Q. What is the change in The market supply function is P 10+Q and the market consumer surplus associated with a minimum floor price of Points out of 1.00 $30 P Flag question Select one: a. $30 b. -$100 C. -$55 d. Zero
Answer #1

Solution -

Option D. Zero is Correct option.

Explanation -

we know that equilibrium condition

QS=QD

10 + Q = 70 - 2Q

3Q = 70 -10

Q=60/3

Q =20

Put Q=20 in QS for Equilibrium Price

P = 10 + Q

P = 10 + 20

P = 30

Equilibrium Price $30 and Price Floor also $30 both are same

it means

there is no change in price and Y axis so

hence proved there is no change in Consumer Surplus

So Option D. Zero is Correct option.